OpenStudy (anonymous):

The first four terms of a sequence are shown below. 7, 4, 1, -2 Which of the following functions best defines this sequence? f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) + 4; for n ≥ 1

OpenStudy (amistre64):

is the next term 3 more, 3 less, 4 more, or 4 less ... than the one before it?

OpenStudy (anonymous):

Could you rephrase that?

OpenStudy (amistre64):

im sure I could ...

OpenStudy (amistre64):

7+n = 4, solve for n

OpenStudy (anonymous):

-3

OpenStudy (amistre64):

then id pick the one that opts for a -3

OpenStudy (anonymous):

So, I'd choose f(1) = 7, f(n + 1) = f(n) - 3; for n ≥ 1 ?

OpenStudy (anonymous):

Thank you, it was correct.